OpenStudy (anonymous):

HELP PLEASE!! MEDALS WILL BE GIVEN!! Solve the system of equations: 9a+3b+c=2 a+b+c=4 6a+b=0

OpenStudy (anonymous):

Pick one equation and solve for one variable. It doesn't matter which one. Show me your work and I will give you the next step.

OpenStudy (anonymous):

b=-6a

OpenStudy (anonymous):

Great. Now substitute -6a for b in one of the other equations.

OpenStudy (anonymous):

9a + 3(-6a) +c =2

OpenStudy (anonymous):

so that equals to -9a+c=2

OpenStudy (anonymous):

Now substitute the same -6a for b in the other 3 variable equation. That should leave you with 2 equations and 2 variables.

OpenStudy (anonymous):

a + -6a +c =4 -5a +c =4

OpenStudy (anonymous):

so then c= 4 +5a

OpenStudy (anonymous):

Don't solve for c on this one. Your two equations are: -9a+c=2 -5a+c=4 Can you solve a systems of equations with two variables? I would suggest subtraction....

OpenStudy (anonymous):

alright thank you :) i got it from here :)

OpenStudy (anonymous):

Great. If you need an additional resource, www.purplemath.com is a great resource for high school math.